let y= cosx+cos2x+..........cos7x divided by sinx+sin2x+ …....sin7x
the value of y when x= pi /8 is.....

Grade:10

2 Answers

Rinkoo Gupta
9 years ago
cospi/8+cos2pi/8+cos3pi/8+cos4pi/8+cos5pi/8+cos6pi/8+cos7pi/8
=cospi/8+cos2pi/8+cos3pi/8+cospi/2+cos(pi-3pi/8)+cps(pi-2pi/8)+cos(pi-pi/8)
=cospi/8+cos2pi/8+cos3pi/8+cospi/2-cos3pi/8-cos2pi/8-cospi/8
=cospi/2
=0
y= (cospi/8+cos2pi/8+cos3pi/8+cos4pi/8+cos5pi/8+cos6pi/8+cos7pi/8)/(sinpi/8+sin2pi/8+sin3pi/8+sin4pi/8+sin5pi/8+sin6pi/8+sin7pi/8
=0/sinpi/8+sin2pi/8+sin3pi/8+sin4pi/8+sin5pi/8+sin6pi/8+sin7pi/8
=0
Thanks & Regards
Rinkoo Gupta
AskIITians Faculty
Arun
9 years ago
Its cos(2*pi/8+(7-1)*pi/8)/sin(2*pi/8+(7-1)*pi/8)=cos(pi)/sin(pi)=-1/0= minus infinity